# Learning to Rank Critical Road Segments via Heterogeneous Graphs with OD Flow Integration

### v2

@@ 4.2 Road segment embedding
In the heterogeneous sequences generated by HetGWalk, road segments, OD pairs, paths, and their associated attributes are interleaved in a way that often induces long-range contextual dependencies. Traditional recurrent architectures such as GRUs (gru) or LSTMs (lstm) rely on sequential inductive biases and therefore struggle to model interactions between elements that are far apart in the sequence, especially when heterogeneous entities participate in multi-hop, cross-type relationships. To effectively capture such long-range contextual dependencies, we adopt a Transformer encoder (transformer), whose fully connected self-attention mechanism directly models all pairwise interactions within each sequence without relying on positional proximity.
Before feeding a sequence into the encoder, each node is linearly projected into a type-specific embedding space, ensuring that heterogeneous entities (OD pairs, paths, segments, and attributes) are represented within a unified latent domaing. Multiple self-attention layers progressively integrate contextual information, producing context-dependent embeddings for each element. For a road segment $v_{i}^{l}$ appearing in multiple sequences, the encoder generates multiple embeddings reflecting its varying context. These embeddings are aggregated using an Attention-based Multiple Instance Learning (AMIL) mechanism (AMIL), yielding a final segment representation $h^{i}$ that adaptively weights each contextual embedding according to its importance.
@@
The multi-head self-attention mechanism in the Transformer captures co-occurrence patterns and functional relationships among road segments from multiple perspectives. When segments frequently co-occur within the same sequences, different attention heads specialize in recognizing distinct types of dependencies: some focus on spatial interactions induced by shared OD flows or functionally related paths, while others concentrate on attribute-driven similarities—such as traffic volume, lane count, and speed category—to identify semantic affinities. This specialized, parallel processing enables segments that serve OD pairs with comparable characteristics (e.g., similar demand levels or matching OD types) to form semantically coherent clusters in the embedding space. As a result, the final embedding for each segment effectively integrates its topological context, OD-driven long-range spatial dependencies, and diverse attribute-driven similarities, yielding a context-aware representation suitable for assessing segment criticality.
@@ 4.3 Listwise learning to rank module
To optimize this module, we adopt the well-established ListNet framework (listnet), specifically its efficient Top-One probability variant, as our training objective. This choice is motivated by its superior alignment with our goal of identifying the most critical segments. The loss function is defined as the Kullback-Leibler (KL) divergence between the ground-truth and predicted probability distributions, which are obtained by applying the SoftMax function to the ground-truth importance scores and the model’s predicted scores, respectively:
@@ 4.4.1 Complexity Analysis of HetGL2R
The computational cost of HetGL2R primarily stems from three core components:
@@
    HetGWalk Sampling Module: The overall computational complexity of the joint random-walk procedure consists of a one-time preprocessing stage and the subsequent walk-generation stage. In preprocessing, Alias tables are constructed for all nodes in the Trip Graph (TG) and the Attribute-Guided Graphs (AGs), incurring a total cost of $O(m+\sum_{\phi\in\{o,p,l\}}a^{\phi})$, where $m$ is the number of TG edges and $a^{\phi}$ is the number of attribute edges in $AG_{\phi}$; The memory requirement is of the same asymptotic order. After preprocessing, each TG step requires one Alias lookup, and each AG step (entity–attribute–entity) requires two, all in $O(1)$ time. Thus, generating numnum walks of length lenlen from each of the $n$ nodes costs $O(num\cdot n\cdot len)$ in total. Overall, the joint random-walk procedure requires $O(m+\sum_{\phi\in\{o,p,l\}}a^{\phi}+num\cdot n\cdot len)$, where the first two terms correspond to preprocessing and the last term accounts for walk generation.
@@
    Transformer Encoder Module: The sequences generated by HetGWalk are fed into a Transformer encoder to learn context-aware node embeddings. For sequences of length lenlen with hidden dimension $d$, the self-attention computation has complexity $O(len^{2}\cdot d)$ per sequence, and the feed-forward network contributes $O(len^{2}\cdot d)$. If numnum walks are generated for each of the $n$ nodes, the total per-epoch cost is $O(num\cdot n\cdot len^{2}\cdot d+num\cdot n\cdot len\cdot d^{2})$, with memory requirements of $O(len^{2})$ for attention maps and $O(len\cdot d)$ for embeddings per sequence.
@@
In summary, the overall complexity of HetGL2R is linear with respect to the number of sampled sequences and quadratic with respect to the sequence length. Its highly parallelizable design (especially in the walk sampling and sequence encoding stages) facilitates efficient training and inference on modern hardware (e.g., GPUs).
@@ 4.4.2 Complexity Comparison with Baseline Methods
    Message-Passing GNNs (e.g., HGT (rf48), GAT (rf50)). These models learn node representations through iterative neighborhood aggregation. Each layer incurs a computational cost of $O(|E|d)$, where $|E|$ denotes the number of edges and d the feature dimensionality. Modeling long-range dependencies requires stacking multiple layers, causing the total complexity to increase with both the graph size and the network depth. In addition, the layer-wise dependency inherent to message passing introduces sequential execution constraints, which fundamentally limit the degree of parallelism achievable within a single batch.
@@
    Random Walk-based Homogeneous Methods (e.g., DeepWalk (rf39)): The sampling stage of these methods shares a similar complexity with our HetGWalk, i.e., $O(num\cdot n\cdot len)$. The key computational difference emerges in the embedding learning phase. These methods typically employ a separate Skip-gram model, which, while highly efficient per training sample due to its shallow architecture and negative sampling, operates on a corpus of generated sequences in a distinct, non-end-to-end step. In contrast, our Transformer encoder, despite having a higher per-sequence complexity of $O(len^{2}\cdot d)$, is integrated into an end-to-end framework. This design eliminates the need for a separate training pipeline and allows for the direct co-optimization of all components, although this integration introduces a higher computational cost per model update.
@@ 5.1 Dataset Construction
To evaluate the proposed framework, three traffic network datasets were constructed using the SUMO (rf2) microscopic traffic simulator. Two datasets, SY-Net110 and SY-Net514, were generated from subregions of the real Shenyang urban road network, containing 110 and 514 nodes, respectively. These networks were extracted from OpenStreetMap and calibrated with realistic road capacities and speed limits. In addition, SY-Net110 and SY-Net514 incorporated real traffic flow data collected during peak hours, from which the corresponding OD flows were inferred (20 OD pairs for SY-Net110 and 90 OD pairs for SY-Net514). Furthermore, a large-scale randomly generated road network (RD-Net3478) with 3,478 nodes and 800 OD pairs was created to assess the model’s scalability to a larger network with a distinct structural topology. For all datasets, SUMO was used to simulate steady-state traffic flows, and the resulting OD–path–segment mappings were extracted to construct heterogeneous trip graphs for model training and evaluation. Table 2 summarizes the statistical properties of the three networks.
To generate the ground-truth ranking of road segment importance, we used the microscopic traffic simulation platform SUMO to simulate traffic dynamics on three networks. For each road segment, we simulated a failure scenario by reducing its maximum speed limit to 10% of its original value. The number of other segments that subsequently experienced failure (defined as a speed drop below 10% of the speed limit) was recorded over multiple time windows. The final importance score (IS (rf8)) of each segment was computed as the weighted sum of affected segments across time, capturing the overall cascading impact of its failure on the network.
@@ 5.2 Baselines and evaluation metrics
To evaluate the effectiveness of the proposed embedding module for ranking tasks, we compared it with the following LTR baseline methods:
  • 1.
    DeepWalk (rf39): Generates node sequences through truncated random walks and learns low-dimensional node embeddings via the Skip-gram model, preserving graph structural proximity in the latent space.
  • 2.
    GraphRNA (rf40): Efficiently integrates topological structures and node attributes through attribute-aware random walks combined with a recurrent neural architecture, enabling context-sensitive graph representation learning.
  • 3.
    TraNode2vec (rf7): Evaluates node importance by combining clustering-based topology analysis with embedding learning, effectively capturing both local and global network structures.
  • 4.
    MGL2Rank (rf8): A random walk–based attributed graph representation learning framework that employs pairwise ranking objectives to model relative node importance relationships.
  • 5.
    HGT (rf48): A heterogeneous graph transformer that models large-scale, type-rich graphs using type-dependent attention mechanisms and a scalable mini-batch sampling strategy, allowing effective learning across diverse node and edge types.
  • 6.
    Dir-GNN (rf53): Improves GNN performance on heterophilic graphs by separately aggregating incoming and outgoing edges in directed graphs while maintaining comparable accuracy on homophilic structures.
  • 7.
    GAT (rf50): Learns node representations through masked self-attention mechanisms that assign adaptive importance weights to neighboring nodes without requiring costly matrix operations.
  • 8.
    ENGINE (rf51): Generates expressive node embeddings by aggregating multi-layer hidden states from pre-trained large language models (LLMs), enabling rich semantic representation for graph-based tasks.
  • 9.
    GNNLLMEmb (rf52): Initializes node features using the last hidden layer representations from an LLM and subsequently applies a GCN to learn topology-enhanced node embeddings that combine linguistic and structural information.
  • 10.
    GATv2 (gatv2): A dynamic graph attention network that modifies the computation order of attention to make neighbor weights explicitly dependent on the query node. It achieves strictly stronger expressiveness and performance than GAT while maintaining a comparable number of parameters.
  • 11.
    PDFormer (PDFormer): A multi-view graph representation model equipped with long-range spatiotemporal attention mechanisms. It simultaneously captures local neighborhood effects and distant coupling relationships, effectively embedding congestion propagation patterns and topological constraints into the node representation space.
  • 12.
    DR-HGNN (DR-HGN): A heterogeneous road network graph model that employs dynamic relation modeling and hierarchical aggregation to jointly encode multiple node and edge types, thereby learning latent representations that integrate both structural and semantic information.
  • 13.
    SeHGNN (SeHGNN): A structure-efficient heterogeneous GNN that precomputes neighbor aggregation to reduce complexity, extends the receptive field through long meta-paths, and leverages Transformer-style semantic fusion to integrate multi-relational semantics, achieving strong representational power with efficient training.
  • 14.
    Edge-GNN (Edge-GNN): An edge-centric message-passing framework that explicitly models edge attributes and connection patterns. It aggregates information from edges to nodes, producing node representations that are more sensitive to traffic-flow propagation.
  • 15.
    LambdaLoss (LambdaLoss): A ranking-oriented loss function that directly optimizes ranking metrics by training a node scoring function to generate importance scores for ordered ranking tasks.
  • 16.
    RankFormer (Rankformer): A Transformer-based listwise learning-to-rank model that introduces listwide labels and jointly optimizes novel listwide objectives with conventional listwise losses, enabling the modeling of inter-item dependencies and improving overall ranking quality.
  • 17.
    LightGBM (lighrgbm): As a gradient boosting–based learning-to-rank baseline, LightGBM is trained with the LambdaMART objective on node attributes and topology-derived features, optimizing overall ranking quality in accordance with the evaluation metrics.
Among these methods, only MGL2Rank LambdaLoss, RankFormer and LightGBM are inherently designed for ranking. The other models are used solely to learn node embeddings, which are then fed into our proposed listwise learning-to-rank module to generate the final ranking results. According to the requirements of large language models, we convert structured or non-textual attributes into textual descriptions as inputs for ENGINE and GNNLLMEmb.
Additionally, we compared HetGL2R with the following non-learnable ranking baseline methods in terms of ranking performance:
  • 1.
    Improved PageRank (IPageRank) (rf3): Combines PageRank centrality and node attributes to rank node importance.
  • 2.
    CRRank (rf1): A heterogeneous graph-based ranking algorithm for road network nodes, considering centrality, attributes, irreplaceability, and upstream-downstream relationships between adjacent nodes.
  • 3.
    OVED-Rank (rf44): Integrating effective distance with multiple topological metrics for efficient evaluation of key spreaders’ influence in networks.
We employ the following metrics to evaluate all methods:
(1) Normalized Discounted Cumulative Gain (NDCG@K): This metric penalizes lower-ranked elements through logarithmic discounting, thereby placing greater emphasis on top-ranked items:
@@
where, $n$ denotes the list length, and $y_{i}$ represents the sorted list.
(2) Wasserstein distance (Earth Mover’s Distance,EMD): Quantifies the discrepancy between the predicted distribution and the true distribution.
(3) Ideal ranking list deviation (Diff) (rf1): Measures the relative ranking error of elements at each position in the list:
@@
(4) Kendall’s rank correlation coefficient (Kendall’s $\tau$) (tau): This metric measures the ordinal association between two rankings by comparing the number of concordant and discordant pairs:
@@
For HetGWalk, we set the propagation factor $e$ to 2, the walk bias factor $\alpha$ to 0.6, the number of sampled sequences to 25, and the sequence length to 20. The embedding dimension of the encoding network is set to 64. The experiments are conducted on the same synthetic dataset under identical experimental settings to ensure fair comparisons.
5.3 Experimental results
For each network (SY-Net110, SY-Net514, and RD-Net3478), 70% of the nodes were randomly selected for training, and the remaining 30% were reserved for testing. Tables 3 - 5 summarize the ranking performance of all methods across multiple evaluation metrics. The results consistently show that learning-based ranking models outperform non-learning baselines across all datasets, while HetGL2R achieves the best overall performance. Specifically, HetGL2R improves $\operatorname{NDCG@K}$, Diff, and $\tau$ by at least 1.82%, 7.52%, and 1.63% on SY-Net110, by 2.98% and 4.40% on $\operatorname{NDCG@K}$ and Diff on SY-Net514 (ranking second only to HGT in $\tau$, with a marginal 0.70% gap), and by 3.39%, 4.98%, and 1.12% on SY-Net3478. These improvements confirm that the proposed embedding module effectively captures both the structural and functional dependencies of road segments, leading to robust and accurate ranking results.
@@ 5.3 Experimental results
Figure 4: The Performance of Each Method on SY-Net110.

### v3

@@ 4.2 Road segment embedding
In the heterogeneous sequences generated by HetGWalk, road segments, OD pairs, paths, and their associated attributes are interleaved in a way that can induces long-range contextual dependencies. Traditional recurrent architectures such as GRUs (Cho et al., 2014) or LSTMs (Graves, 2012) rely on sequential inductive biases and may be less effective in modeling interactions between elements that are far apart in the sequence, especially when heterogeneous entities participate in multi-hop, cross-type relationships. To capture such long-range contextual dependencies, we adopt a Transformer encoder (Vaswani et al., 2017), whose self-attention mechanism enables direct modeling of pairwise interactions within each sequence without relying on positional proximity.
Before feeding a sequence into the encoder, each node is linearly projected into a type-specific embedding space, ensuring that heterogeneous entities (OD pairs, paths, segments, and attributes) are represented within a unified latent domaing. Multiple self-attention layers progressively integrate contextual information, producing context-dependent embeddings for each element. For a road segment $v_{i}^{l}$ appearing in multiple sequences, the encoder generates multiple embeddings reflecting its varying context. These embeddings are aggregated using an attention-based multiple instance learning (AMIL) mechanism (Ilse et al., 2018), yielding a final segment representation $h^{i}$ that adaptively weights each contextual embedding according to its relative importance.
@@
The multi-head self-attention mechanism allows the model to capture co-occurrence patterns and functional relationships among road segments from multiple perspectives. When segments frequently co-occur within the same sequences, different attention heads can capture diverse types of dependencies, including spatial interactions associated with shared OD flows or related paths, as well as attribute-based similarities such as traffic volume, lane count, and speed category. This parallel processing facilitates the formation of semantically coherent segment representations in the embedding space. As a result, the final embedding for each segment integrates topological context, OD-induced long-range dependencies, and attribute-based similarities, yielding a context-aware representation suitable for assessing segment criticality.
@@ 4.3 Listwise learning to rank module
To optimize this module, we adopt the well-established ListNet framework (Cao et al., 2007), specifically its efficient Top-One probability variant, as our training objective. This choice is motivated by its superior alignment with our goal of identifying the most critical segments. The loss function is defined as the Kullback-Leibler (KL) divergence between the ground-truth and predicted probability distributions, which are obtained by applying the SoftMax function to the ground-truth importance scores and the model’s predicted scores, respectively:
@@ 4.4.1 Complexity Analysis of HetGL2R
The construction of the heterogeneous trip graph is performed once in an offline manner and scales linearly with the number of OD pairs, paths, and road segments, without requiring iterative optimization. Consequently, the overall computational cost of HetGL2R is dominated by the following three core components:
@@
    HetGWalk Sampling Module: The overall computational complexity of the joint random-walk procedure consists of a one-time preprocessing stage and the subsequent walk-generation stage. In preprocessing, Alias tables are constructed for all nodes in the Trip Graph (TG) and the Attribute-Guided Graphs (AGs), incurring a total cost of $O(m+\sum_{\phi\in\{o,p,l\}}a^{\phi})$, where $m$ is the number of TG edges and $a^{\phi}$ is the number of attribute edges in $AG_{\phi}$; The memory requirement is of the same asymptotic order. After preprocessing, each TG step requires one Alias lookup, and each AG step (entity–attribute–entity) requires two, all in $O(1)$ time. Thus, generating numnum walks of length lenlen from each of the $n$ nodes costs $O(num\cdot n\cdot len)$ in total. Overall, the joint random-walk procedure requires $O(m+\sum_{\phi\in\{o,p,l\}}a^{\phi}+num\cdot n\cdot len)$, where the first two terms correspond to preprocessing and the last term accounts for walk generation. All heterogeneous random walks are generated with fixed length and bounded sampling budgets, and rely only on local transitions. As a result, the walk generation process scales linearly with the number of sampled walks and can be efficiently parallelized.
@@
    Transformer Encoder Module: The sequences generated by HetGWalk are fed into a Transformer encoder to learn context-aware node embeddings. For sequences of length lenlen with hidden dimension $d$, the self-attention computation has complexity $O(len^{2}\cdot d)$ per sequence, and the feed-forward network contributes $O(len^{2}\cdot d)$. If numnum walks are generated for each of the $n$ nodes, the total per-epoch cost is $O(num\cdot n\cdot len^{2}\cdot d+num\cdot n\cdot len\cdot d^{2})$, with memory requirements of $O(len^{2})$ for attention maps and $O(len\cdot d)$ for embeddings per sequence. Notably, the Transformer encoder is applied to sampled walk sequences rather than the entire heterogeneous graph. Therefore, its computational cost depends on the sequence length instead of the global network size, avoiding the quadratic complexity associated with full-graph attention.
@@
From an implementation perspective, the main components of HetGL2R exhibit inherent parallelism. Random walk sampling and sequence encoding are performed on independent samples, while the listwise ranking stage introduces minimal additional overhead. As a result, the overall complexity of HetGL2R is linear with respect to the number of sampled sequences and quadratic with respect to the sequence length, supporting efficient training and inference in large-scale settings using standard parallel computing resources.
@@ 4.4.2 Complexity Comparison with Baseline Methods
    Message-Passing GNNs (e.g., HGT (Hu et al., 2020), GAT (Velickovic et al., 2018)). These models learn node representations through iterative neighborhood aggregation. Each layer incurs a computational cost of $O(|E|d)$, where $|E|$ denotes the number of edges and d the feature dimensionality. Modeling long-range dependencies requires stacking multiple layers, causing the total complexity to increase with both the graph size and the network depth. In addition, the layer-wise dependency inherent to message passing introduces sequential execution constraints, which fundamentally limit the degree of parallelism achievable within a single batch.
@@
    Random Walk-based Homogeneous Methods (e.g., DeepWalk (Perozzi et al., 2014)): The sampling stage of these methods shares a similar complexity with our HetGWalk, i.e., $O(num\cdot n\cdot len)$. The key computational difference emerges in the embedding learning phase. These methods typically employ a separate Skip-gram model, which, while highly efficient per training sample due to its shallow architecture and negative sampling, operates on a corpus of generated sequences in a distinct, non-end-to-end step. In contrast, our Transformer encoder, despite having a higher per-sequence complexity of $O(len^{2}\cdot d)$, is integrated into an end-to-end framework. This design eliminates the need for a separate training pipeline and allows for the direct co-optimization of all components, although this integration introduces a higher computational cost per model update.
@@ 5.1 Dataset Construction
To evaluate the proposed framework, three traffic network datasets were constructed using the SUMO (Lopez et al., 2018) microscopic traffic simulator. Two datasets, SY-Net110 and SY-Net514, were generated from subregions of the real Shenyang urban road network, containing 110 and 514 nodes, respectively. These networks were extracted from OpenStreetMap and calibrated with realistic road capacities and speed limits. Moreover, SY-Net110 and SY-Net514 incorporated real peak-hour traffic flow data, from which the corresponding OD flows were inferred (20 OD pairs for SY-Net110 and 90 OD pairs for SY-Net514). In addition, a large-scale randomly generated road network, RD-Net3478, with 3,478 nodes and 800 OD pairs, was constructed to evaluate the scalability of the proposed model on a network with a distinct structural topology. Table 2 summarizes the statistical properties of the three networks.
To obtain a ground-truth ranking of road segment importance, we conducted simulation-based failure experiments using SUMO by reducing the maximum speed limit of each segment to 10% of its original value. The number of other segments that subsequently experienced failure—operationally defined as a speed drop below 10% of the corresponding speed limit—was recorded over multiple time windows. The final importance score (IS (Xu and Zhang, 2024)) for each segment was computed as a weighted sum of the affected segments across time, capturing the cascading impact of a segment failure on overall network performance.
@@ 5.2 Baselines and evaluation metrics
To evaluate the effectiveness of the proposed method, we compare it with a broad set of baselines, which we group into two main categories: learning-based methods and rule-based methods.
The learning-based baselines are further divided into two types:
(1) Representation learning (embedding) models that focus on learning node representations from graph topology, attributes, or language model prompts. These include DeepWalk (Perozzi et al., 2014), GraphRNA (Huang et al., 2019), TraNode2vec (Huang et al., 2022), HGT (Hu et al., 2020), Dir-GNN (Rossi et al., 2024), GAT (Velickovic et al., 2018), GATv2 (Brody et al., 2021), PDFormer (Jiang et al., 2023), DR-HGNN (Liu et al., 2025), SeHGNN (Yang et al., 2023), Edge-GNN (Jana et al., 2023), ENGINE (Zhu et al., 2024), and GNNLLMEmb (Wu et al., 2025). For ENGINE and GNNLLMEmb, structured attributes are converted into textual descriptions to meet language model input requirements.
(2) Specialized learning-to-rank (LTR) models that are explicitly designed for ranking tasks. These include MGL2Rank (Xu and Zhang, 2024), LambdaLoss (Wang et al., 2018), RankFormer (Buyl et al., 2023), and LightGBM (Ke et al., 2017).
The rule-based baselines rank nodes based on topology- and attribute-driven heuristics without learnable parameters. These include Improved PageRank (IPageRank) (Agryzkov et al., 2012), CRRank (Xu et al., 2018), and OVED-Rank (Ullah et al., 2025).
To ensure a fair comparison, the node embeddings produced by all models in group (1) are fed into our proposed listwise LTR module to generate the final rankings.
@@
Note: Best results are highlighted in bold. Higher values indicate better performance for NDCG@K and Kendall’s $\tau$, while lower values are better for EMD and Diff.
We employ the following metrics to evaluate all methods:
(1) Normalized Discounted Cumulative Gain (NDCG@K) evaluates ranking quality by assigning higher weights to top-ranked elements through logarithmic discounting, thereby emphasizing the correctness of the most important items:
@@
where, $n$ denotes the number of nodes, and $y_{i}$ represents the sorted list.
(2) Wasserstein distance (Earth Mover’s Distance, EMD) quantifies the discrepancy between the predicted and ground-truth importance distributions, capturing differences in both ranking order and magnitude.
(3) Ideal ranking list deviation (Diff) (Xu et al., 2018) measures the relative deviation between the predicted ranking and the ideal ranking at each list position, reflecting position-sensitive ranking errors:
@@
(4) Kendall’s rank correlation coefficient (Kendall’s $\tau$) (KENDALL, 1938) assesses the ordinal consistency between two rankings by comparing the numbers of concordant and discordant element pairs:
@@
For HetGWalk, we set the propagation factor $e$ to 2, the walk bias factor $\alpha$ to 0.6, the number of sampled sequences to 25, and the sequence length to 20. The encoder consists of 6 layers, each with 8 attention heads and an embedding dimension of 64. We optimize the model using the Adam optimizer with a learning rate of 0.001. All experiments are conducted on the same synthetic dataset under these settings to ensure a fair comparison.
Figure 4: Performance comparison of different methods on SY-Net110. The horizontal axis denotes the predicted rank of road segments, while the vertical axis represents their ground-truth importance scores (IS). Each point corresponds to a road segment, and different colors indicate different baseline methods and the proposed HetGL2R.
@@
5.3 Experimental results
For each network (SY-Net110, SY-Net514, and RD-Net3478), 70% of the nodes were randomly selected for training, with the remaining 30% reserved for testing. Tables 3 - 5 report the ranking performance of all methods across multiple evaluation metrics. The results consistently show that learning-based ranking models outperform non-learning baselines across all datasets, while HetGL2R achieves the best overall performance. Specifically, HetGL2R improves $\operatorname{NDCG@K}$, Diff, and $\tau$ by at least 1.82%, 7.52%, and 1.63% on SY-Net110, by 2.98% and 4.40% on $\operatorname{NDCG@K}$ and Diff on SY-Net514 (ranking second only to HGT in $\tau$, with a marginal 0.70% gap), and by 3.39%, 4.98%, and 1.12% on SY-Net3478. These improvements confirm that the proposed embedding module effectively captures both the structural and functional dependencies of road segments, leading to robust and accurate ranking results.